in a throw of two dice,the probability throwing doublets not more than twice is

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Required probabilty=1-6/36=1-1/6=5/6

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Aman Bansal
13 years ago

dear pankaj,

in a single throw of two dice the probabiity of throwing doublets not more than twicw is 1/6

sample spaces=36

favourable outcomes =6

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL